OpenStudy (anonymous):

Evaluate integral: integral (3^(-2xln3))

OpenStudy (anonymous):

\[\int3^{-2x\ln3}~dx\] Let \(u=-2x\ln3\), so that \(-\dfrac{1}{2\ln3}du=dx\), or \(-\dfrac{1}{2\ln3}du=dx\). \[-\frac{1}{2\ln3}\int3^u~du\] Now use some exponent/log properties: \[3^u=e^{\ln3^u}=e^{u\ln3}\] \[-\frac{1}{2\ln3}\int e^{u\ln3}~du\] Let \(t=u\ln3\), so that \(\dfrac{1}{\ln3}dt=du\): \[-\frac{1}{2\ln3}\cdot\frac{1}{\ln3}\int e^t~dt\] \[-\frac{1}{2(\ln3)^2}\int e^t~dt\]